The perimeter of a triangle is 24 feet. The second side is two feet longer thatn the first. The third side is two feet longer than the second. What are the lengths of the sides?
Answer by nerdybill(7384)

The perimeter of a triangle is 24 feet. The second side is two feet longer than the first. The third side is two feet longer than the second. What are the lengths of the sides?
.
Let x = length of first side
then from "The second side is two feet longer than the first."
x+2 = length of second side
and from "The third side is two feet longer than the second."
x+2+2 = x+4 = length of third side
.
x + x+2 + x+4 = 24
3x+6 = 24
3x = 18
x = 6 feet (first side)
.
second side:
x+2 = 6+2 = 8 feet
.
third side:
x+4 = 6+4 = 10 feet
.
Answer: 6,8 and 10 feet